Question 1 (10 points)
Find the distance between the two points (3,-4) and (-4,5)
Question 2 (10 points)
The idea of eccentricity is associated to which one of the four types of conic sections?
Question 3 (5 points)
Each of the next four questions refer to the parabola y = 5(x+2)^2 + 7.
Find the vertex of the parabola

Question 4 (5 points)
Find the focus of the parabola
Question 5 (5 points)
The directrix line of the parabola is the line y =
Question 6 (5 points)
The length of the latus rectum of this parabola is
Question 7 (20 points)
Find the standard form of the circle x^2 + y^2 + 2x - 6y - 6 = 0.
Question 8 (5 points)
For the circle x^2 + y^2 + 2x - 6y - 6 = 0 of the previous problem, find the center.
Question 9 (5 points)
For the same circle x^2 + y^2 + 2x - 6y - 6 = 0 of the previous two questions, find the radius.
Question 10 (20 points)
Find the standard form of the ellipse given by the equation x^2 + 25y^2 -2x + 150y + 201 = 0.
Question 11 (5 points)
Here is the equation of an ellipse in standard form: (1/16)(x + 2)^2 + (1/9)(y - 5)^2 = 1 . Each of the next four questions refer to this ellipse. Find the center of the ellipse.

Question 12 (5 points)
Find the foci of the ellipse.
Question 13 (5 points)
Find the length of the major axis.
Question 14 (5 points)
Find the distance between the two foci.
Question 15 (20 points)
Find the standard form of the hyperbola given by the equation 4x2 – 25y2 – 50y – 125 = 0.
Question 16 (5 points)
Here is the equation of a hyperbola in standard form: (1/10)(x - 1)^2 - (y - 1)2 = 1 . Each of the next four questions refer to this hyperbola. Find the center of the hyperbola.

Question 17 (5 points)
Find the foci of the hyperbola.
Question 18 (5 points)
One of the two axes (transverse or conjugate) is parallel to the y-axis. Find its length.
Question 19 (5 points)
Find the two asymptotes for the hyperbola

Answer by checkley71(8403)

QUESTION1) THE DISTANCE OF TWO OF THE SIDES ARE 5-(-4)=9 & 3-(-4)=7.
NOW THAT WE HAVE TWO SIDES WE SOVE FOR THE HYPOTENUSE (C) IN THE FORMULA
A^2=B^2=C^2
9*9+7*7=C^2
81+49=C^2
C^2=130
C=SQRT130
C=11.401754
----------------------------
HERE'S THE SOLUTION TO YOUR FIRST PROBLEM, HOWEVER:
WE ARE HERE TO HELP SOLVE INDIVIDUAL PROBLEMS BUT NOT TO DO ALL YOUR HOMEWORK.